OpenStudy (anonymous):

PLEEEAAAASSSEE HELP!!!!!!!!!!!!!! Y varies inversely as three times x. When x=6, y=2. Find x when y=12. Choices: x=1 x=3 x=6 x=9 PLEASE HELP

OpenStudy (dls):

y propotional to 1/x so... \[\Huge y = \frac{k}{x}\] Try now..find k from given values and solve accordingly.

OpenStudy (anonymous):

for k I got 9

OpenStudy (anonymous):

@DLS

OpenStudy (dls):

x=6 then y=2 \[\Huge 2=\frac{k}{6},k=?\] not sure how you got 9

OpenStudy (anonymous):

It says y varies inversely as three times x. So wouldn't you need to multiply x by 3? So the equation would be 2=k/6(3) and when you simplify it you would get 2=k/18 and to find k you would need to divide 18/2 which will give you k which is 9

OpenStudy (anonymous):

@DLS

OpenStudy (dls):

then k=36